OpenStudy (anonymous):

How do you solve this? 2x + 4y - 3z = -7

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

only one equation? O.o

OpenStudy (anonymous):

Lol no, there are 3 of them

OpenStudy (anonymous):

oh okay lol, can you post them too?

OpenStudy (anonymous):

2x + 4y - 3z = -7 3x + y + 4z = -12 x + 3y + 4z = 4

OpenStudy (anonymous):

alright :) 1. Let's label them first 2x + 4y - 3z = -7 --> eqn 1 3x + y + 4z = -12 --> eqn 2 x + 3y + 4z = 4 --> eqn 3 2. Choose which method you want to use, elimination or substitution? Elimination is the best & easiest way. Do you know how to apply this method?

OpenStudy (anonymous):

I haven't done it in a while, I just did substitution in my last equation, but haven't tried elimination.

OpenStudy (anonymous):

ok. For solving 3 variable equations, we have to pick two equations first. we will use the 1st and 2nd equation. 2x + 4y - 3z = -7 --> eqn 1 3x + y + 4z = -12 --> eqn 2 Let's say we want to eliminate the variable y, what will you multiply to eqn 2 so that when you add eqn 1 and eqn 2, y will get cancelled?

OpenStudy (anonymous):

-4?

OpenStudy (anonymous):

right! now multiply -4 for each term in eqn 2 -4(3x + y + 4z = -12) -4(3x) + -4(y) + -4(z) = -4(-12) when you simplify this, what will you get?

OpenStudy (anonymous):

-12x - 4y -4z = 48 ?

OpenStudy (anonymous):

perfect :P now add that to eqn 1 2x + 4y - 3z = -7 --> eqn 1 -12x - 4y -4z = 48 --> eqn (2 multiplied by -4) -----------------

OpenStudy (anonymous):

I am subtracting those, right?

OpenStudy (anonymous):

since you already multiplied 2nd eqn with -4, you have to add them I'll do the 1st term for you :) 2x + 4y - 3z = -7 --> eqn 1 -12x - 4y -4z = 48 --> eqn (2 multiplied by -4) ----------------- -10 x then continue it, y will get cancelled. Then add the terms with z and so on

OpenStudy (anonymous):

-10x -7z = 41 ?

OpenStudy (anonymous):

yes. -10x - 7z = 41 let this be your eqn 4 Now, next step is you have to pick another pair from your equations that include eqn 3 and do this process again. Remember that we are eliminating variable y okay? :)

OpenStudy (anonymous):

Just to be sure, now I take -10x - 7z = 41 and x + 3y + 4z = 4 and do everything we just did?

OpenStudy (anonymous):

No, cause it doesn't have a y anymore

OpenStudy (anonymous):

no, just leave eqn 4 (which is -10x - 7z = 41) for now. We will use it later. Just pick any pairs: eqn 1 and eqn 3 or eqn 2 and eqn 3? After you choose, do the process again and eliminate y

OpenStudy (anonymous):

Okay, so if I take eqn 1 and eqn 3 how would I eliminate the y is it's a 4 and a 3?

OpenStudy (anonymous):

Sorry if I am asking basic questions, math is not my strong suit

OpenStudy (anonymous):

No worries! we're here for a reason :) 2x + 4y - 3z = -7 --> eqn 1 x + 3y + 4z = 4 --> eqn 3 just think about the least common factor of 4 and 3, which is 12. If you make those y-terms 12 and -12, they will get cancelled right? So for the 1st eqn, what will you multiply 4 to get 12? For eqn 3, what will you multiply 3, to get -12?

OpenStudy (anonymous):

1st eqn by 3 and 3rd eqn by -4

OpenStudy (anonymous):

right :) So multiply eqn 1 by 3 and multiply eqn 3 by -4 ... then add these equations together

OpenStudy (anonymous):

I got 3x - 7z = -3 ?

OpenStudy (anonymous):

Oh i forgot to multiply it to everything in the equation!

OpenStudy (anonymous):

let me see :) 3(2x + 4y - 3z = -7) --> eqn 1 x3 6x + 12y -9z = -21 -4(x + 3y + 4z = 4 ) --> eqn 3x -4 -4x - 12y -16z = -16 6x + 12y -9z = -21 -4x - 12y -16z = -16 ------------------ 2x - 25z = -37 hmm I get a different one. :) It's okay, so now let this be eqn 5 Now let's combine eqn 4 by eqn 5: -10x - 7z = 41 --> eqn 4 2x - 25z = -37 --> eqn 5 You can use either substitution or elimination to combine these two and solve for the variables x and z. Can you manage this part?

OpenStudy (anonymous):

Okay, I was going to do substitution but then I got confused, where I would put 41 and -37?

OpenStudy (anonymous):

I'm not sure what you mean... can you show me your steps so that I can check? :)

OpenStudy (anonymous):

Well I'm not really sure what to do, as this lesson is about learning sub and elimination, but I think I would start by putting something like -10x -7 ( ? ) = ? do you get what i'm saying?

OpenStudy (anonymous):

yes, that's right. -10x - 7( ? ) = 41 You have to rearrange eqn in terms of z so that you can substitute it into the eqn 4. 2x - 25z = -37 this should be z=

OpenStudy (anonymous):

what do you mean? z = ?

OpenStudy (anonymous):

for example: you have equation 2x - 3z =4 arrange this in terms of z so it will be like